OpenStudy (anonymous):

if f(x)=6x and g(x)+4x+1, find (f0g)(x). a]24x+6 b]24x^2+6x c]10x+1 d]24x+1

OpenStudy (anonymous):

between the f and g its not a letter nor a number its a circle that's not shaded in in the middle

jigglypuff314 (jigglypuff314):

(fog)x = f(g(x)) plug in 4x + 1 for g(x) f(4x+1) then plug in (4x+1) for x in 6x so 6(4x+1) = ?

OpenStudy (anonymous):

im not sure :(

jigglypuff314 (jigglypuff314):

distributive property is like a(b+c) = ab + ac so for example 5(2x+3) = 5(2x) + 5(3) = 10x + 15 knowing that, what is 6(4x+1) = ?

OpenStudy (anonymous):

oh 24x+6?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

so the answer would be a?

OpenStudy (anonymous):

omg thank you so much!:)

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

wait I have one more question if you don't mind

jigglypuff314 (jigglypuff314):

sure :)